OpenStudy (naveenbhatia1312):

Compute the exact value of the function for the given x-value without using a calculator. f(x)=(1/5)^x for x=-1

OpenStudy (misty1212):

HI!!

OpenStudy (naveenbhatia1312):

Hey

OpenStudy (misty1212):

on no \[\huge f(x)=\left(\frac{1}{5}\right)^{x}\]right?

OpenStudy (naveenbhatia1312):

yes thats the equation

OpenStudy (misty1212):

then \[\huge f(-1)=\left(\frac{1}{5}\right)^{-1}\]

OpenStudy (misty1212):

the minus one in the exponent means "take the reciprocal" i.e flip it

OpenStudy (naveenbhatia1312):

so just poitive 5

OpenStudy (misty1212):

yup

OpenStudy (naveenbhatia1312):

thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]